OpenStudy (anonymous):

A class consists of 3 boys and 6 girls willing to form 3 groups of 3 called Groups A, B, C. How many ways can such distinct groups be formed?

OpenStudy (anonymous):

@hartnn @ganeshie8

OpenStudy (anonymous):

@thomaster

OpenStudy (anonymous):

@myininaya

OpenStudy (anonymous):

So boys and girls are not distinct? LOL!

OpenStudy (anonymous):

My strategy is: if you assign the girls to each group, then there is only one way to assign the remaining boys.

OpenStudy (anonymous):

And vise versa. In fact in this case it looks like it would be easier to assign the boys into each group.

OpenStudy (anonymous):

Distinct as in they have different names and the groups are noted. Meaning, order is important.

OpenStudy (anonymous):

@wio Can you give a mathematical interpretation?

OpenStudy (anonymous):

Seems like there is no individuality here. There is only distinction by gender.

OpenStudy (anonymous):

You want to know how I'd assign the boys to the groups?

OpenStudy (anonymous):

@wio The rest of the question is actually cut off. This is just the first part. The other parts talk about assigning boys and girls separately. For now, I just want to know how to do this part.

OpenStudy (anonymous):

Okay let me get this straight: The kids are distinct. The gender doesn't matter then.

OpenStudy (anonymous):

You have \(3+6=9\) total kids.

OpenStudy (anonymous):

Gender doesn't matter yet. Not in this part of the question. Yes, total is 9.

OpenStudy (anonymous):

The ways to divide \(9\) kids among \(3\) groups of size \(3\) is \[ \frac{9!}{3!3!3!} \] http://www.wolframalpha.com/input/?i=9!%2F(3!3!3!)

OpenStudy (anonymous):

Does that take into account that the groups have different names? Shouldn't one multiply by 3! because of the ways they can choose names?

OpenStudy (anonymous):

Or you could do it this way: \[ \binom 93\times \binom 63 \times \binom 33 \] http://www.wolframalpha.com/input/?i=(9+choose+3)(6+choose+3)(3+choose+3)

OpenStudy (anonymous):

That's what I did at first. I understand how that works. I'm just concerned about the fact that they can take separate names afterward.

OpenStudy (anonymous):

So you want it to be: 1a) choose a group 1b) choose the group members 2a) ... ?

OpenStudy (anonymous):

Like this? \[ \binom 31\binom 93\times \binom 21\binom 63 \times \binom 11\binom 33 \] http://www.wolframalpha.com/input/?i=%283+choose+1%29%289+choose+3%29+%282+choose+1%29%286+choose+3%29+%281+choose+1%29%283+choose+3%29

OpenStudy (anonymous):

Okay, which one is the proper way?

OpenStudy (anonymous):

Exact question says "How many ways are there to assign 9 of them to Groups A, B C?

OpenStudy (anonymous):

I suppose the second way then.

OpenStudy (anonymous):

Why not just *3! ? Can you explain what you did there?

OpenStudy (anonymous):

@wio ?

OpenStudy (anonymous):

I chose one group before choosing it's members

OpenStudy (anonymous):

How about: Divide them into 3 groups Assign the groups names A,B,C. *3! ?

OpenStudy (anonymous):

You can do that too.

OpenStudy (anonymous):

Okay, so in what case is the first methid accurate? The first one you did?

OpenStudy (anonymous):

*method

OpenStudy (anonymous):

Umm, the first method is accurate if the groups are indistinguishable, I believe.

OpenStudy (anonymous):

Okay, thanks. @wio How many ways are there to assign students to these groups such that every group includes 1 boy?

OpenStudy (anonymous):

1a) choose a group 1b) choose a boy 1c) choose two girs

OpenStudy (anonymous):

\[ \binom 31 \binom 31 \binom 62 \times \binom 21 \binom 21 \binom 42 \times \binom 11 \binom 11 \binom 22 \]

OpenStudy (anonymous):

Are you sure? Seems quite large. I was thinking: \[\frac{ 6! }{ 2!2!2! } *3!\]

OpenStudy (anonymous):

Thought: Partition girls into three groups of 2. Then multiply by the number of ways the boys can fill the slots.

OpenStudy (anonymous):

Are the groups no longer distinct?

OpenStudy (anonymous):

They always are,

OpenStudy (anonymous):

@Mertsj @agent0smith What do you think?

OpenStudy (anonymous):

It's getting confusing because of the fact that the groups have names.

OpenStudy (mertsj):

I am not good at probability but I can tell you what I think. For the first group we have( 9 x 8 x 7)/3! which is 84 For the second group we have (6x5x4)/3! which is 20 For the last group we have (3x2x1)?3! which is 1 The grand total is 105

OpenStudy (anonymous):

Trust my answer. =) I already explained how it works fundamentally.

OpenStudy (agent0smith):

I think wio's answers here are on-point.

OpenStudy (anonymous):

Unless you can find where I'm double counting or something.

OpenStudy (agent0smith):

\[\large \binom 31\binom 93\times \binom 21\binom 63 \times \binom 11\binom 33\] and \[\large \binom 31 \binom 31 \binom 62 \times \binom 21 \binom 21 \binom 42 \times \binom 11 \binom 11 \binom 22\]